# Deep Ignorance: Filtering Pretraining Data Builds Tamper-Resistant Safeguards into Open-Weight LLMs

### v1

@@ 1 Introduction
As frontier large language models (LLMs) grow more advanced, their developers have raised concerns about increasing security risks posed by their models. For example, in its Gemini 2.5 Pro Preview model card, Google Deepmind remarks, “subsequent revisions in the next few months could lead to a model that reaches the critical capability level [for harmful novice uplift](Google AI / DeepMind Gemini Team, 2025). Anthropic reported that it was preemptively activating its Safety Level 3 protocols for Claude Opus 4, writing that they “could not rule out” “the ability to significantly assist individuals or groups with basic STEM backgrounds in obtaining, producing, or deploying CBRN [^3] Chemical, Biological, Radiological, and Nuclear (CBRN) weapons.(Anthropic, 2025). OpenAI’s ChatGPT Agent system card states “We have decided to [precautionarily] treat this launch as high capability in the Biological and Chemical domain(OpenAI, 2025b).
Today, frontier LLMs such as the above are often deployed with closed weights behind APIs. However, open-weight models are being released at an increasing rate (Bhandari et al., 2025), and their capabilities tend to lag only six to twelve months behind those of closed-weight models (Cottier et al., 2024; Maslej et al., 2024). Open models offer unique benefits related to transparency, research, and the deconcentration of power (Bommasani et al., 2024; Kapoor et al., 2024; François et al., 2025). However, they create unique risks from downstream modifications and unmonitored use (Seger et al., 2023; Chan et al., 2023; Huang et al., 2024). This prompts the question: How can harmful uses of advanced open-weight models be effectively mitigated?
Despite the rising prominence of open-weight models, the science of open-weight model safety is nascent (François et al., 2025). Closed-weight deployments keep a model’s weights secure and allow for monitors and filters on its inputs and outputs. However, the defining challenge for open-weight model safety is that open models can be modified arbitrarily by downstream actors. Despite recent work on developing tamper-resistant models (e.g. Henderson et al., 2023; Rosati et al., 2024; Tamirisa et al., ; Sheshadri et al., 2024), existing techniques can consistently be undone within several hundred fine-tuning steps or fewer (e.g. Qi et al., 2024b; Huang et al., 2025; Che et al., 2025; Fan et al., 2025; Hu et al., 2024; Sheshadri et al., 2024; Łucki et al., 2024; Qian et al., 2025; Deeb & Roger, 2024).
@@
    Defense in Depth: We demonstrate that data filtering cannot prevent LLMs from leveraging harmful knowledge provided in-context, but that Circuit-Breaking-based techniques (Zou et al., 2024) offer complementary defenses. However, we show that none of the defenses we test are resistant to staged attacks that combine fine-tuning and in-context retrieval (Section 4).
@@ 2.3 Baseline Post-Training Safeguards
We used versions of CB (Zou et al., 2024) with and without targeted LAT (Sheshadri et al., 2024). [^14] We opted not to use TAR (Tamirisa et al., ) due to its computationally expensive and delicate nature. Prior work from Che et al. (2025) and Qi et al. (2024b) found that TAR both struggled to resist fine-tuning attacks and suffered from significant dysfluency and off-target capability degradation. We also note that CB+LAT (Zou et al., 2024; Sheshadri et al., 2024) is algorithmically similar to TAR, with the key difference being the parameterization of the adversarial attacker.
@@ 2.4 Defining and Measuring Biothreat Proxy Knowledge
We focus on biology knowledge that can serve as a proxy for biorisk-relevant information. To measure this knowledge, we utilize WMDP-Bio (Li et al., 2024b), a public four-way multiple-choice question answering (MCQA) benchmark developed by subject matter experts. This benchmark was designed to assess biothreat-related yet harmless knowledge of dual-use biological processes and laboratory techniques. It is widely used to benchmark LLM safeguards (e.g., Thaker et al., 2024; Deeb & Roger, 2024; Łucki et al., 2024; Kolbeinsson et al., 2024). The WMDP-Bio eval contains questions about proxy information from six key topics. Here, we list these six topics alongside representative example questions.
@@
The WMDP-Bio MCQA benchmark does not evaluate LLM’s biorisk capabilities directly, which may require other advanced capabilities such as multimodal reasoning and agentic tool-use. It instead evaluates factual knowledge that related to biorisk topics. Other works have evaluated biorisk-related knowledge using challenging multimodal MCQA evaluations (Götting et al., 2025) and with red-teaming exercises (Mouton et al., 2023). While MCQA formats have known limitations, they provide a standardized and scalable way to assess LLMs’ internalized knowledge and are a practical evaluation format for our base models (Biderman et al., 2024).
@@ 3.1 Model Tampering Attacks
  • Latent-space attacks: Following Sheshadri et al. (2024), we develop latent-space prompt perturbations at layers 0, 8, 16, 24, and 30. We optimized these attacks to be universal, making the model respond correctly on a set of 32 held-out questions. Through hyperparameter tuning, we chose to use 64 optimization steps with a batch size of 16 using gradient descent (not projected gradient descent), a learning rate of $10^{-3}$, and exponential learning rate decay with $\gamma=0.95$. The prompt whose tokens we applied the latent-space attacks to depended on the evaluation prompt used for MCQA vs. Cloze evaluations. We run each attack 3 times and report the mean and standard deviation in plots.
@@
  • Benign fine-tuning: Safeguards should ideally persist through legitimate model adaptation. For example, organizations routinely fine-tune models on internal datasets for domain-specific tasks. However, prior work has found that non-adversarial tampering can also undo LLM safeguards (Qi et al., 2023; Deeb & Roger, 2024; Che et al., 2025). This presents an exceptionally low-cost attack since attackers do not need access to domain-specific data. To test models against benign tampering, we fine-tune them identically to our adversarial fine-tuning attacks except on the WikiText dataset (Merity et al., 2016) instead of the WMDP-Bio Forget set.
@@ 3.2 Filtering Resists Fine-Tuning Attacks for up to 10k Steps and 300M Tokens
LLM safeguards are known to be vulnerable to benign interventions (e.g., (Qi et al., 2023; Che et al., 2025; Deeb & Roger, 2024; Pandey et al., 2025). Figure 5 confirms this for CB and CB+LAT. However, our filtered models’ biothreat-proxy performance remains unchanged under benign fine-tuning on WikiText (Merity et al., 2016).
@@
Fine-tuning attack configurations are difficult to compare, but in Appendix E, Table 10 we compare reported details from prior works on adversarial fine-tuning attacks on open-weight language models $\geq$ 1B parameters in size (Fan et al., 2025; Hu et al., 2024; Sheshadri et al., 2024; Che et al., 2025; Łucki et al., 2024; Qian et al., 2025; Deeb & Roger, 2024; Muhamed et al., 2025; Tamirisa et al., ; Qi et al., 2023). We conduct fine-tuning attacks for the largest number of unique examples, total steps, and (step $\times$ batch size) product compared to any of these related works.
Figure 6: Pretraining data filtering cannot prevent in-context retrieval of unwanted information, but Circuit-Breaking can. However, no models resist an ensemble fine-tuning+in-context-retrieval attack. Baseline and filtered models alike can perform well on our “open-book” biothreat knowledge tests in which a passage containing the answer is given in context. Circuit-Breaking complements filtering by impairing the model’s ability to retrieve biothreat-related information in-context. No defenses, however, resist our ensemble attack.
@@
4 Defense in Depth
In safety-critical settings, using multiple layers of redundant safeguards is a common strategy to minimize risks. As a result, defense-in-depth (also known as the Swiss Cheese model) is a ubiquitous principle in safety engineering (Reason, 1990). Meanwhile, prior works on defending AI systems against attacks have found that combinations of defenses often yield useful synergies (Raff et al., 2019; Kolbeinsson et al., 2024; McKenzie et al., 2025).
@@ 4.1 Data Filtering and Circuit-Breaking are Complementary
LLMs do not necessarily need to know harmful information to provide it. For example, “AI Agents” are increasingly popular and often include tools that allow the base language model to search and retrieve information from the web (Gao et al., 2023; Xi et al., 2025; Casper et al., 2025). Meanwhile, Yu et al. (2025) showed that these types of agents may pose acute safety challenges because LLMs often fail to behave safely when augmented with retrieval tools. As a result, in cases when harmful information is easily searchable, retrieval augmentation may hinder the effectiveness of training data filtering. [^15] We note that in cases when search-tool augmentation attacks are possible, LLMs are likely to offer limited uplift because such a search tool could be used by humans directly without the LLM.
@@
To test model safety under retrieval augmentation, we designed a 1,000-question multiple-choice biothreat proxy knowledge benchmark with two evaluation modes: “closed-book” and “open-book.” Using 1,000 biology paper abstracts from the WMDP-Bio Forget set, we prompted Claude 3.7 Sonnet (Anthropic, 2025a) to generate multiple choice questions about biothreats that (1) had a single correct answer that would be known to experts irrespective of context, (2) were hard to answer by nonexperts without context, and (3) were trivial for nonexperts to answer with the paper’s abstract as context. See the prompt that we used in Section I.1. Figure 6 shows that our filtered models perform well on the open-book evaluation, while our filtered models with CB and CB+LAT resist correctly answering biology questions even when the answer is given in context. However, we also find that no models are resistant to an ensemble fine-tuning + open-book attack, suggesting that such ensemble attacks may be a weak point of this type of defense-in-depth strategy. [^16] While a limitation in the open-weight context, the ability to leverage dual-use information in context can be a useful property of data filtering in a closed-weight context. For instance, since dual-use knowledge by definition can have benign applications, model providers could permit their LLMs to access dual-use knowledge databases only when interacting with trusted users. Providers could then restrict these knowledge databases for untrusted users. This setup would continue to allow LLMs access to dual-use knowledge for prosocial outcomes.
@@ 5 Related Work
Filtering harmful contents from training data is recognized as a key risk management technique (e.g., Korbak et al., 2023; Thorn, 2025; François et al., 2025). However, curating web-scale datasets is difficult (Paullada et al., 2021). Aside from the high, direct costs it incurs (Ngo et al., 2021), it also suffers from filtering errors (Ziegler et al., 2022), degradation of dataset quality (Welbl et al., 2021), the massively multilingual nature of internet text (Kreutzer et al., 2022), cultural biases in content moderation (Welbl et al., 2021; Dodge et al., 2021; Xu et al., 2021; Stranisci & Hardmeier, 2025), and the inherently contextual nature of harmfulness. Modern data corpora that are used to train LLMs have been found to contain harmful, toxic, and abusive content (Birhane et al., 2023b; a; Thiel, 2023). Concurrently with this project, some frontier model developers have publicly mentioned efforts to filter pretraining data for harmful content (Kamath et al., 2025; OpenAI, 2025a; Meta, 2025; OpenAI, 2025; Anthropic, 2025c; OpenAI, 2025). None of which, however, provides precise details about what data was filtered, how it was filtered, how much was filtered, or how the success of the filtering was evaluated. On the other hand, several concurrent works have openly studied pretraining data filtering and its effects on safety in language models at the $\leq$ 2B parameter scale. Maini et al. (2025) and Li et al. (2025a), both studied filtering for harmful and toxic content. Closely related to our work, Lee et al. (2025) used model distillation on filtered data in a proof of concept for robust capability suppression by training on curated data, however, they only experimented with models up to 500M parameters and fine-tuning attacks up to 500 steps of adversarial fine-tuning. In Section 3.2, we show competitive and robust capability suppression via pretraining data filtering at the 6.9B scale and test on up to 10,000 steps and 305M tokens of adversarial fine-tuning.
@@
Aside from fine-tuning LLMs to refuse harmful requests (e.g., (Mazeika et al., 2024; Liu et al., 2024; Yu et al., 2024; Casper et al., 2024; Sheshadri et al., 2024; Zou et al., 2024; Tamirisa et al., )), some ‘machine unlearning’ techniques have also been studied for their ability to directly suppress harmful LLM capabilities (Barez et al., 2025; Liu et al., 2025a). Many of these techniques have been proposed (Li et al., 2025b; Liu et al., 2025b) including fine-tuning-based methods (e.g., (Jang et al., 2022; Eldan & Russinovich, 2023; Li et al., 2024b; Zou et al., 2024; Sheshadri et al., 2024; Tamirisa et al., ; Gandikota et al., 2024; Rosati et al., 2024; Qian et al., 2025; Anthropic Alignment Team, 2025)) and mechanistic-interventions (e.g., (Muhamed et al., 2025; Lo et al., 2024; Guo et al., 2024; Wang et al., 2025; Michaud et al., 2025; Schoepf et al., 2025)). However, despite recent efforts, existing approaches to machine unlearning are vulnerable to attacks (Łucki et al., 2024; Che et al., 2025); prone to major side-effects (Qi et al., 2024b); difficult to evaluate (Feng et al., 2025); and generally fraught with conceptual, technical, and practical challenges (Cooper et al., 2024).
@@
Much recent work has also focused on eliciting harmful capabilities from LLMs using white-box techniques. These include embedding-space (soft prompt) attacks (Schwinn et al., 2023; 2024; Yang et al., 2024; Geisler et al., 2024; Xhonneux et al., 2024), latent-space attacks (Casper et al., 2024; Sheshadri et al., 2024; Fort, 2023; Kirch et al., 2024; Zou et al., 2023a; Wang & Shu, 2023), and weight-space (fine-tuning) attacks on both benign and adversarial data (Jain et al., 2023; Yang et al., 2023; Qi et al., 2023; Bhardwaj & Poria, 2023; Lermen et al., 2023; Zhan et al., 2023; Wei et al., 2024; Ji et al., 2024; Qi et al., 2024a; Hu et al., 2024; Halawi et al., 2024; Greenblatt et al., 2024; Li et al., 2024c; Hofstätter et al., 2025; Deeb & Roger, 2024; Qi et al., 2024b; Che et al., 2025; Tamirisa et al., ; Fan et al., 2025; Łucki et al., 2024; Hsiung et al., 2025; Pandey et al., 2025; Wallace et al., 2025). Past works have found that few-shot fine-tuning attacks are particularly effective at efficiently eliciting latent capabilities from LLMs (Greenblatt et al., 2024; Qi et al., 2024b; Huang et al., 2024; Hofstätter et al., 2025; Che et al., 2025) with state-of-the-art tamper resistance methods only conferring resistance up to dozens or hundreds of examples of fine-tuning (Sheshadri et al., 2024; Qi et al., 2024b; Łucki et al., 2024; Huang et al., 2025; Qian et al., 2025; Che et al., 2025). In Section 3, we achieve state-of-the-art tamper robustness through filtering pre-training data.

### v2

@@ 1 Introduction
As frontier large language models (LLMs) grow more advanced, their developers have raised concerns about increasing security risks posed by their models. For example, in its Gemini 2.5 Pro Preview model card, Google Deepmind remarks, “subsequent revisions in the next few months could lead to a model that reaches the critical capability level [for harmful novice uplift](Google AI / DeepMind Gemini Team, 2025). Anthropic reported that it was preemptively activating its Safety Level 3 protocols for Claude Opus 4, writing that they “could not rule out” “the ability to significantly assist individuals or groups with basic STEM backgrounds in obtaining, producing, or deploying CBRN [^3] Chemical, Biological, Radiological, and Nuclear (CBRN) weapons.(Anthropic, 2025). OpenAI’s ChatGPT Agent system card states “We have decided to [precautionarily] treat this launch as high capability in the Biological and Chemical domain(OpenAI, 2025a).
Today, frontier LLMs such as the above are often deployed with closed weights behind APIs. However, open-weight models are being released at an increasing rate (Bhandari et al., 2025), and their capabilities tend to lag only six to twelve months behind those of closed-weight models (Cottier et al., 2024; Maslej et al., 2024). Open models offer unique benefits related to transparency, research, and the deconcentration of power (Bommasani et al., 2024; Kapoor et al., 2024; Eiras et al., 2024; françois2025differentapproachaisafety). However, they create unique risks from downstream modifications and unmonitored use (Seger et al., 2023; Chan et al., 2023; Huang et al., 2024; Eiras et al., 2024). This prompts the question: How can harmful uses of advanced open-weight models be effectively mitigated?
Despite the rising prominence of open-weight models, the science of open-weight model safety is nascent (françois2025differentapproachaisafety; Srikumar et al., 2024). Closed-weight deployments keep a model’s weights secure and allow for monitors and filters on its inputs and outputs. However, the defining challenge for open-weight model safety is that open models can be modified arbitrarily by downstream actors. Despite recent work on developing tamper-resistant models (e.g. Henderson et al., 2023; Rosati et al., 2024; Tamirisa et al., ; Sheshadri et al., 2024), existing techniques can consistently be undone within several hundred fine-tuning steps or fewer (e.g. Qi et al., 2024b; Huang et al., 2025; Che et al., 2025; Fan et al., 2025; Hu et al., 2024; Sheshadri et al., 2024; Łucki et al., 2024; Qian et al., 2025; Deeb and Roger, 2024).
@@
    Defense-in-Depth: We demonstrate that data filtering cannot prevent LLMs from leveraging harmful knowledge provided in-context, but that Circuit-Breaking-based techniques (Zou et al., 2024) offer complementary defenses. However, we show that none of the defenses we test are resistant to staged attacks that combine fine-tuning and in-context retrieval (Section 4).
@@ 2.3 Baseline Post-Training Safeguards
Prior benchmarking work from Che et al. (2025) found that CB (Zou et al., 2024) was state-of-the-art for tamper-resistance. Meanwhile, (Sheshadri et al., 2024) found that latent adversarial training could be applied to further improve tamper resistance of post-training algorithms. See also Table 10 for a more in-depth discussion of prior works’ testing of tamper-resistant safeguards. [^14] We opted not to use TAR (Tamirisa et al., ) due to its computationally expensive and delicate nature. Prior work from Che et al. (2025) and Qi et al. (2024b) found that TAR both struggled to resist fine-tuning attacks and suffered from significant dysfluency and off-target capability degradation. We also note that CB+LAT (Zou et al., 2024; Sheshadri et al., 2024) is algorithmically similar to TAR, with the key difference being the parameterization of the adversarial attacker.
@@ 2.4 Defining and Measuring Biothreat Proxy Knowledge
We focus on biology knowledge that can serve as a proxy for biorisk-relevant information. To measure this knowledge, we utilize WMDP-Bio (Li et al., 2024b), a public four-way multiple-choice question answering (MCQA) benchmark developed by subject matter experts. This benchmark was designed to assess biothreat-related yet harmless knowledge of dual-use biological processes and laboratory techniques. It is widely used to benchmark LLM safeguards (e.g., Thaker et al., 2024; Deeb and Roger, 2024; Łucki et al., 2024; Kolbeinsson et al., 2024). The WMDP-Bio eval contains questions about proxy information from six key topics. Here, we list these six topics alongside representative example questions.
@@
The WMDP-Bio MCQA benchmark does not evaluate LLM’s biorisk capabilities directly, which may require other advanced capabilities such as multimodal reasoning and agentic tool-use. It instead evaluates factual knowledge that related to biorisk topics. Other works have evaluated biorisk-related knowledge using challenging multimodal MCQA evaluations (Götting et al., 2025) and with red-teaming exercises (Mouton et al., 2023). While MCQA formats have known limitations (Khatun and Brown, 2024), they provide a standardized and scalable way to assess LLMs’ internalized knowledge and are a practical evaluation format for our base models (Biderman et al., 2024).
@@ 3.1 Model Tampering Attacks
  • Latent-space attacks: Following Che et al. (2025), we develop latent-space prompt perturbations at layers 0, 8, 16, 24, and 30. We optimized these attacks to be universal, making the model respond correctly on a set of 32 held-out questions. Through hyperparameter tuning, we chose to use 64 optimization steps with a batch size of 16 using gradient descent (not projected gradient descent), a learning rate of $10^{-3}$, and exponential learning rate decay with $\gamma=0.95$. The prompt whose tokens we applied the latent-space attacks to depended on the evaluation prompt used for MCQA vs. Cloze evaluations. We run each attack 3 times and report the mean and standard deviation in plots.
@@
  • Benign fine-tuning: Safeguards should ideally persist through legitimate model adaptation. For example, organizations routinely fine-tune models on internal datasets for domain-specific tasks. However, prior work has found that non-adversarial tampering can also undo LLM safeguards (Qi et al., 2023; Deeb and Roger, 2024; Che et al., 2025). This presents an exceptionally low-cost attack since attackers do not need access to domain-specific data. To test models against benign tampering, we fine-tune them identically to our adversarial fine-tuning attacks except on the WikiText dataset (Merity et al., 2016) instead of the WMDP-Bio Forget set.
@@ 3.2 Filtering Resists Fine-Tuning Attacks for up to 10k Steps and 300M Tokens
LLM safeguards are known to be vulnerable to benign interventions (e.g., (Qi et al., 2023; Che et al., 2025; Deeb and Roger, 2024; Pandey et al., 2025). Figure 5 confirms this for CB and CB+LAT. However, our filtered models’ biothreat-proxy performance remains unchanged under benign fine-tuning on WikiText (Merity et al., 2016).
@@
Fine-tuning attack configurations are difficult to compare, but in Appendix E, Table 10 we compare reported details from prior works on adversarial fine-tuning attacks on open-weight language models $\geq$ 1B parameters in size (Fan et al., 2025; Hu et al., 2024; Sheshadri et al., 2024; Che et al., 2025; Łucki et al., 2024; Qian et al., 2025; Deeb and Roger, 2024; Muhamed et al., 2025; Tamirisa et al., ; Qi et al., 2023). We conduct fine-tuning attacks for the largest number of unique examples, total steps, and (step $\times$ batch size) product compared to any of these related works.
Figure 6: Pretraining data filtering cannot prevent in-context retrieval of unwanted information, but Circuit-Breaking can. However, no models resist an ensemble fine-tuning + in-context-retrieval attack. Baseline and filtered models alike can perform well on our “open-book” biothreat knowledge tests in which a passage containing the answer is given in context. Circuit-Breaking complements filtering by impairing the model’s ability to retrieve biothreat-related information in context. No defenses, however, resist our ensemble attack.
@@
4 Defense-in-Depth
Is data curation all you need? In safety-critical settings, using multiple layers of redundant safeguards is a common strategy to minimize risks. As a result, defense-in-depth (also known as the Swiss Cheese model) is a ubiquitous principle in safety engineering (Reason, 1990). Meanwhile, prior works on defending AI systems against attacks have found that combinations of defenses often yield useful synergies (Raff et al., 2019; Kolbeinsson et al., 2024; McKenzie et al., 2025).
@@ 4.1 Data Filtering and Circuit-Breaking are Complementary
LLMs do not necessarily need to know harmful information to provide it (Shumailov et al., 2024). For example, “AI Agents” are increasingly popular and often include tools that allow the base language model to search and retrieve information from the web (Gao et al., 2023; Xi et al., 2025; Casper et al., 2025). Meanwhile, Yu et al. (2025) showed that these types of agents may pose acute safety challenges because LLMs often fail to behave safely when augmented with retrieval tools. As a result, in cases when harmful information is easily searchable, retrieval augmentation may hinder the effectiveness of training data filtering. [^15] We note that in cases when search-tool augmentation attacks are possible, LLMs are likely to offer limited uplift because such a search tool could be used by humans directly without the LLM.
@@
To test model safety under retrieval augmentation, we designed a 1,000-question multiple-choice biothreat proxy knowledge benchmark with two evaluation modes: “closed-book” and “open-book.” Using 1,000 biology paper abstracts from the WMDP-Bio Forget set, we prompted Claude 3.7 Sonnet (Anthropic, 2025a) to generate multiple choice questions about biothreats that (1) had a single correct answer that would be known to experts irrespective of context, (2) were hard to answer by nonexperts without context, and (3) were trivial for nonexperts to answer with the paper’s abstract as context. See the prompt that we used in Section I.1. Figure 6 shows that our filtered models perform well on the open-book evaluation, while our filtered models with CB and CB+LAT resist correctly answering biology questions even when the answer is given in context. This shows how filtering and CB can be complementary. However, we also find that no models are resistant to an ensemble fine-tuning + open-book attack, suggesting that such ensemble attacks may be a weak point of this type of defense-in-depth strategy. [^16] While a limitation in the open-weight context, the ability to leverage dual-use information in context can be a useful property of data filtering in a closed-weight context. For instance, since dual-use knowledge by definition can have benign applications, model providers could permit their LLMs to access dual-use knowledge databases only when interacting with trusted users. Providers could then restrict these knowledge databases for untrusted users. This setup would continue to allow LLMs access to dual-use knowledge for prosocial outcomes.
@@ 5 Related Work
Filtering harmful contents from training data is recognized as a key risk management technique (e.g., Korbak et al., 2023; Thorn, 2025; françois2025differentapproachaisafety; Srikumar et al., 2024). However, curating web-scale datasets is difficult (Paullada et al., 2021). Aside from the high, direct costs it incurs (Ngo et al., 2021), it also suffers from filtering errors (Ziegler et al., 2022), degradation of dataset quality (Welbl et al., 2021), the massively multilingual nature of internet text (Kreutzer et al., 2022), cultural biases in content moderation (Welbl et al., 2021; Dodge et al., 2021; Xu et al., 2021; Stranisci and Hardmeier, 2025), and the inherently contextual nature of harmfulness. Modern data corpora that are used to train LLMs have been found to contain harmful, toxic, and abusive content (Birhane et al., 2023b; a; Thiel, 2023). Concurrently with this project, some frontier model developers have publicly mentioned efforts to filter pretraining data for harmful content (Kamath et al., 2025; OpenAI, 2025b; Meta, 2025; OpenAI, 2025d; Anthropic, 2025c; OpenAI, 2025c). None of which, however, provides precise details about what data was filtered, how it was filtered, how much was filtered, or how the success of the filtering was evaluated. On the other hand, several concurrent works have openly studied pretraining data filtering and its effects on safety in language models at the $\leq$ 2B parameter scale. Maini et al. (2025) and Li et al. (2025a), both studied filtering for harmful and toxic content. Closely related to our work, Lee et al. (2025) used model distillation on filtered data in a proof of concept for robust capability suppression by training on curated data, however, they only experimented with models up to 500M parameters and fine-tuning attacks up to 500 steps of adversarial fine-tuning. In Section 3.2, we show competitive and robust capability suppression via pretraining data filtering at the 6.9B scale and test on up to 10,000 steps and 305M tokens of adversarial fine-tuning.
@@
Aside from fine-tuning LLMs to refuse harmful requests (e.g., (Mazeika et al., 2024; Liu et al., 2024; Yu et al., 2024; Casper et al., 2024; Sheshadri et al., 2024; Zou et al., 2024; Tamirisa et al., )), some ‘machine unlearning’ techniques have also been studied for their ability to directly suppress harmful LLM capabilities (Barez et al., 2025; Liu et al., 2025a). Many of these techniques have been proposed (Li et al., 2025b; Liu et al., 2025b) including fine-tuning-based methods (e.g., (Jang et al., 2022; Eldan and Russinovich, 2023; Li et al., 2024b; Zou et al., 2024; Sheshadri et al., 2024; Tamirisa et al., ; Gandikota et al., 2024; Rosati et al., 2024; Qian et al., 2025; Anthropic Alignment Team, 2025)) and mechanistic-interventions (e.g., (Muhamed et al., 2025; Lo et al., 2024; Guo et al., 2024; Wang et al., 2025; Michaud et al., 2025; Schoepf et al., 2025)). However, despite recent efforts, existing approaches to machine unlearning are vulnerable to attacks (Łucki et al., 2024; Che et al., 2025); prone to major side-effects (Qi et al., 2024b); difficult to evaluate (Feng et al., 2025); and generally fraught with conceptual, technical, and practical challenges (Cooper et al., 2024).
@@
Much recent work has also focused on eliciting harmful capabilities from LLMs using white-box techniques. These include embedding-space (soft prompt) attacks (Schwinn et al., 2023; 2024; Yang et al., 2024; Geisler et al., 2024; Xhonneux et al., 2024), latent-space attacks (Casper et al., 2024; Sheshadri et al., 2024; Fort, 2023; Kirch et al., 2024; Zou et al., 2023a; Wang and Shu, 2023), and weight-space (fine-tuning) attacks on both benign and adversarial data (Jain et al., 2023; Yang et al., 2023; Qi et al., 2023; Bhardwaj and Poria, 2023; Lermen et al., 2023; Zhan et al., 2023; Wei et al., 2024; Ji et al., 2024; Qi et al., 2024a; Hu et al., 2024; Halawi et al., 2024; Greenblatt et al., 2024; Li et al., 2024c; Hofstätter et al., 2025; Deeb and Roger, 2024; Qi et al., 2024b; Che et al., 2025; Tamirisa et al., ; Fan et al., 2025; Łucki et al., 2024; Hsiung et al., 2025; Pandey et al., 2025; Wallace et al., 2025; Hossain et al., 2025). Past works have found that few-shot fine-tuning attacks are particularly effective at efficiently eliciting latent capabilities from LLMs (Greenblatt et al., 2024; Qi et al., 2024b; Huang et al., 2024; Hofstätter et al., 2025; Che et al., 2025) with state-of-the-art tamper resistance methods only conferring resistance up to dozens or hundreds of examples of fine-tuning (Sheshadri et al., 2024; Qi et al., 2024b; Łucki et al., 2024; Huang et al., 2025; Qian et al., 2025; Che et al., 2025). In Section 3, we achieve state-of-the-art tamper robustness through filtering pre-training data.